# Missing Money and Market-Based Adequacy in Deeply Decarbonized Power Systems with Long-Duration Energy Storage

### v1

@@ Abstract
Index Terms:
Long-Duration Energy Storage, Missing Money Problem, Capacity Market, Capacity Accreditation
@@
$q_{\omega ts}^{ch},q_{\omega ts}^{dis}$
@@
$d_{\omega t}^{fix},d_{\omega t}^{flex}$
@@
$D_{\omega t}^{fix},D_{\omega t}^{flex}$
@@ I Introduction
Recent long‐term planning studies have underscored the indispensable role of large‐scale energy storage (ES) in achieving deep decarbonisation of the power sector [1]. Modelling analyses consistently suggest that in futures with very tight greenhouse gas (GHG) emission targets, the additional energy capacity (MWh) requirement is much higher relative to the additional power capacity (MW), indicating that investments in long duration energy storage (LDES) technologies [^1] While LDES does not have a common definition, it usually refers to a class of storage technologies with discharge duration that exceeds 6 hours [2]. will be essential [3]. This need becomes particularly acute in the full or partial absence of alternative clean, firm resources, such as new nuclear power or natural‐gas plants equipped with carbon capture and storage (CCS) [4].
@@
Integration of LDES faces at least two challenges. First, many LDES technologies are capital-intensive and relatively immature. To guide technology development, a recent research stream explored cost thresholds for LDES to appear in least-cost capacity mixes, concluding that significant reductions in energy-capacity costs are needed to compete with other firm capacity resources [5]. Second, the business models that allow LDES to valorize all its operational and environmental benefits for the system [6] have not yet been established. In energy markets with price caps lower than the value of lost load (VOLL), a business model purely based on energy arbitrage remains uneconomic, with narrow price spreads failing to yield sufficient returns to fully support investments [7].
@@
Pure arbitrage is equally unattractive for short-duration energy storage (SDES), e.g., electrochemical batteries. However, in the case of SDES, stacking revenues from multiple market products such as energy and ancillary services [8], combined with generous subsidies, such as tax credits or grants [9], has boosted batteries’ profits and deployment. These two drivers have created robust SDES business models, leading to a sixfold surge of global annual installations since 2020 [10]. However, these models will likely not be transferable to LDES. Short-term applications, such as frequency regulating reserves, and diurnal arbitrage, will continue to be dominated by SDES and additionally face market saturation risks [11]. Furthermore, subsidies and grants can become unsustainable if a large number of clean energy assets needs support.
@@
One benefit of LDES lies in its peaking potential and its contribution to system adequacy [6]. As such, in decarbonized power systems, LDES primarily replaces dispatchable conventional peakers, limits renewable curtailment, and ensures adequacy during contingencies or prolonged periods with low‑VRE generation (so‑called ’energy droughts’ [12]). Consequently, as with peaking units, a substantial portion of the LDES revenues is likely to be realized during scarcity or near-scarcity events. In markets with price caps lower than the VOLL, insufficiently high price spikes during scarcity events could create a missing money problem [13], challenging the viability of investment in LDES. Furthermore, the concentration of scarcity events within particular weather years can amplify interannual profit variability, undermining investor confidence even when wholesale energy price arbitrage provides sufficient revenue over the financial lifetime of the asset [14].
@@
Capacity remuneration mechanisms (CRMs), such as capacity markets (CMs), address missing money and revenue volatility in price‑capped markets. However, CMs were originally designed with thermal generation in mind, and their effectiveness in contexts with substantial ES capacity, particularly LDES, remains understudied. Recent research has focused on accurately valuing the contribution of energy storage to system adequacy. Zachary et al. [15] demonstrated that a change in expected unserved energy (EUE) is a suitable metric for evaluating the contribution of resources to system adequacy, extending the analyses to VRE and storage (yet with strong assumptions about its dispatch). They also demonstrated that a marginal, rather than an average, approach could yield accurate investment signals [^2] Marginal capacity credit reflects the incremental contribution of a new resource to system adequacy, while average capacity credit captures the overall contribution of all existing resources of that type.. Although marginal accreditation methods are more complex to apply, multiple CM operators, including NYISO [16], ISO-NE [17], PJM [18], and GB [19], have recently begun to progress in this direction.
@@
The shift towards marginal accreditation is promising, as investment decisions are highly sensitive to the chosen methodology and underlying assumptions [20]. Yet, even when marginal credits are used for ES and VRE, there is no general proof ensuring that investments will be efficient. Wang et al. [21] demonstrated through simulations that marginal accreditation in CMs can provide price signals that encourage near-efficient investment for VRE and SDES. While these results are encouraging for the effectiveness of CMs in the near future, with narrow participation of ES, they are of limited use for future systems where LDES may play a dominant role.
@@
In general, most existing studies focus on ES with a single value (usually short) of duration or study systems where ES is a minor contributor to a system’s resource adequacy [21, 22]. Many that include ES in CRMs often make simplified assumptions about accreditation [23]. These simplifications make the results of existing studies less relevant for systems with LDES. The duration of LDES, for many emerging technologies, is a design choice independent of the power capacity. Moreover, LDES duration can range from a few to several hundred hours. Different durations yield distinct contributions to system adequacy, complicating decisions for both investors and regulators.
@@
We contribute to the literature in several ways. Methodologically, we extend the framework of [21] by incorporating a price-elastic demand to address degeneracy and improve price interpretability under high VRE and storage penetration. This is also the first study to endogenise continuous ES accreditation within an investment model, using piecewise linear functions of duration. This improves on previous single-duration methods, which misestimate the adequacy value of storage systems with different durations.
Our analysis offers new insights into market-based adequacy in low-carbon systems with high LDES levels. We confirm that the missing money problem remains significant under price caps, but declines as storage is able to capture wider spreads. We also show that the price cap affects revenues and operations below the cap itself, particularly in cases of low CO2 emissions. Well-calibrated CMs with marginal EUE-based accreditation maintain near-efficient investment signals, but their effectiveness is lower under tighter decarbonization targets.
These findings are relevant for CM design and resource adequacy methodologies. They also inform policymakers considering LDES support mechanisms, such as the UK’s cap-and-floor scheme [24] by highlighting the interplay between accreditation, revenue volatility, and revenue sufficiency.
@@ II-A1 Generators
Let $\mathcal{G}$ be a set of electricity generators. Each generator has three decision variables: its installed capacity ($c_{g}$), the capacity contracted by the CM ($c^{CM}_{g}$), and the electricity produced in each time interval ($q_{\omega,t,g}$), which are in the set $\Gamma^{g}=\left\{q_{\omega,t,g},c_{g},c^{CM}_{g}\geq 0\right\}$. Acting as a price taker, each generator solves the optimization problem (1) below.
$\forall g\in\mathcal{G}\quad\max_{\Gamma^{g}}\sum_{\omega t}w_{\omega t}\delta_{\omega}\left(\lambda^{E}_{\omega,t}-C_{g}^{V}-\lambda^{CT}EF_{g}\right)q_{\omega tg}$
$+c^{CM}_{g}\lambda^{CM}-c_{g}\left(\widetilde{I}_{g}+I_{g}^{F}\right)\hskip 0.0pt$ (1a)
$q_{\omega tg}\leq c_{g}A_{\omega tg}\hskip 0.0pt\quad\forall\omega,t;\quad(\nu^{C}_{\omega tg})$ (1b)
$c^{CM}_{g}\leq CC_{g}c_{g}\hskip 0.0pt$ (1c)
@@ II-A2 Storage
Let set $\mathcal{S}$ include all ES market participants. ES participants determine power and energy capacity independently. Additionally, ES decision variables include the capacity contracted by the CM, its initial state-of-charge (SOC), its charging, and discharging. All decision variables are in $\Gamma_{s}=\left\{q_{\omega,t,s}^{\text{dis}},q_{\omega,t,s}^{ch},c_{s}^{P},c_{s}^{E},e_{\omega,t,s},e^{init},c^{CM}_{s}\geq 0\right\}$ and their values are determined by solving the optimization problem (2).
@@
$+c^{CM}_{s}\lambda^{CM}-c_{s}^{P}\left({\widetilde{I}}_{s}^{P}+I_{s}^{P.F}\right)-c_{s}^{E}\left({\widetilde{I}}_{s}^{E}+I_{s}^{E.F}\right)\hskip 0.0pt$ (2a)
@@
$e_{\omega ts}=e_{s}^{init}+w_{\omega t}(q_{\omega ts}^{\text{ch}}\eta_{s}^{\text{ch}}-\frac{q_{\omega ts}^{\text{dis}}}{\eta_{s}^{\text{dis}}})\quad\forall\;\omega,\;t=1$ (2e)
$e_{\omega ts}\geq e_{s}^{init}\hskip 0.0pt\;\forall\;\omega,\;t=T$ (2f)
@@
$c^{CM}_{s}\leq c_{s}^{P}f^{CC}(c_{s}^{E},c_{s}^{P})\hskip 0.0pt$ (2h)
Eq. (2a) is the ES’s objective function, maximising its profit. Eq. (2b)-(2c) limit charging and discharging based on installed power capacity. Eq. (2d)-(2e) calculate the energy stored at the end of a time interval of duration $w_{\omega,t}$ by considering the energy stored at the start of the interval and the charging, and discharging during the interval. Eq. (2f)-(2g) limit the energy stored relative to the energy stored at the start of the model’s horizon and the installed energy capacity, respectively [^3] Constraining end-of-horizon storage to exceed initial levels may be restrictive, but it provides a practical way to handle model end effects under limited foresight. It is also consistent with EU rules requiring minimum gas storage levels before winter [25]..
Finally, (2h) limits the capacity contracted by the CM, $c^{CM}_{s}$, to the qualifying capacity of ES. We determine the capacity credit of ES in a novel manner as a function, $f^{CC}$, of ES duration $\zeta_{s}$ $\eta_{s}^{\text{dis}}c_{s}^{E}/c_{s}^{P}$. This new way allows us to properly consider the capacity market contributions of a range of ES durations. The function $f^{CC}$ is piece-wise linear with $|L|$ segments, where each segment $l$ covers a range of durations $\zeta_{s,l-1},\zeta_{s,l}]$. As such, $f^{CC}$ is represented as
$f^{CC}(\zeta_{s})=\alpha_{ls}+\beta_{ls}\zeta_{s},\quad\forall\;\zeta_{s}\in[\zeta_{sl-1},\zeta_{sl}],l\in\mathcal{L}$ (3)
Capacity credit curves are monotonically increasing and concave with respect to duration [26]. In other words, the marginal contribution to adequacy diminishes as duration increases (i.e., the value of $\beta_{l,s}$ is positive but lower for higher durations). We estimate the coefficients of $f^{CC}$, i.e., $\alpha_{l,s},\beta_{l,s}$, using piece-wise linear regression.
In the maximisation problem (2), constraint (2h) can be included as is because $f^{CC}$ is concave. If $f^{CC}$ was not concave, binary variables and associated constraints would be needed to ensure that the segments of the piecewise curve are activated in a logical order. Multiplying both sides of (3) with the installed power capacity $c_{s}^{P}$, (2h) becomes:
$c^{CM}_{s}\leq\alpha_{l,s}c_{s}^{P}+\beta_{l,s}\eta_{s}^{\text{dis}}c_{s}^{E},\quad\forall l\in\mathcal{L}$ (4)
@@ II-A3 Consumers
We assume a single representative agent acting on behalf of all consumers. Its decision variables include the electricity consumption and capacity purchased in the CM, grouped in the set $\Gamma^{C}=\left\{\left.d^{E}_{\omega,t}\right|t\in\mathcal{T},\omega\in\Omega,d^{C}\geq 0\right\}$. Consumers maximise the expected consumer surplus, which is defined as the difference between the integral of willingness to pay (WTP) functions for energy $f^{E}$ and capacity $f^{C}$ and the procurement cost (product of price and quantity purchased).
$\max_{\Gamma^{C}}\sum_{\omega,t}w_{\omega t}\delta_{\omega}\Bigg{[}\int_{0}^{d^{E}_{\omega,t}}f^{E}_{\omega,t}(x)dx-\lambda^{E}_{\omega,t}d^{E}_{\omega,t}\Bigg{]}\\ +\int_{0}^{d^{C}}f^{C}(x)dx-\lambda^{C}d^{C}\hskip 0.0pt$ (5a)
The demand curves for electricity can take diverse shapes, ranging from perfectly price-inelastic to perfectly price-elastic curves. In this work, in line with [27], we assume that the demand curve for electricity at each time interval has two segments. The first segment is perfectly elastic and demands electricity between 0 and $D^{fix}_{\omega,t}$ with a willingness to pay up to $B$. The second segment is a price-elastic linear demand curve where the willingness to pay ranges from 0 to $B$ for consumption between 0 and $D^{flex}_{\omega,t}$. This demand curve helps ensure price uniqueness and resolve dual degeneracy issues in systems with high shares of renewables and storage. The WTP function for energy $f^{E}$ is included in the model using the equations below.
@@
$d_{\omega t}^{\text{flex}}\leq D_{\omega t}^{\text{flex}}\quad\forall\omega t\hskip 0.0pt$ (6c)
The CM function is represented similarly, and it often has more segments with differing levels of elasticity. The CM function $f^{C}(d^{C})$ represents the WTP for capacity, and is a monotonically decreasing function. It consists of $N$ segments, with differing levels of elasticity. The first segment is perfectly price-elastic, willing to pay the capacity market price cap $B^{C}_{1}$ for any capacity between 0 and $D^{\text{C-fix}}$. In the last segment, the WTP $B^{C}_{N}$ decreases to zero as $d^{C}$ reaches the maximum capacity the CM aims to procure. The N2N-2 segments in between are flexible (lengths $D_{n}^{\text{C-flex}}$, starting prices $B^{C}_{n}$, with $B^{C}_{N}=0$) and they are parametrized based on net-CONE. Given this structure, the function can be represented as:
$\int_{0}^{d^{C}=d^{\text{fix}}+\sum_{n=1}^{N-1}d_{n}^{\text{flex}}}f^{C}(x)\,dx=B^{C}_{1}d^{\text{C-fix}}+\\ \sum_{n=1}^{N-1}B^{C}_{n}d_{n}^{\text{C-flex}}-\sum_{n=1}^{N-1}\frac{(d_{n}^{\text{C-flex}})^{2}(B^{C}_{n}-B^{C}_{n+1})}{2D_{n}^{\text{C-flex}}},\hskip 0.0pt$ (7)

### v2

@@ Abstract
$q_{\omega ts}^{\text{ch}},q_{\omega ts}^{\text{dis}}$
@@
$d_{\omega t}^{\text{fix}},d_{\omega t}^{\text{flex}}$
@@
$D_{\omega t}^{\text{fix}},D_{\omega t}^{\text{flex}}$
@@
$\text{VOLL},\text{PC}$
Value of lost load, price cap
@@ I Introduction
Recent long‐term planning studies have underscored the indispensable role of large‐scale energy storage (ES) in achieving deep decarbonisation of the power sector [jafari_power_2020]. Modelling analyses consistently suggest that in futures with very low greenhouse gas (GHG) emission targets, the additional energy capacity (MWh) requirement is much higher relative to the additional power capacity (MW), indicating that investments in long duration energy storage (LDES) technologies [^1] While LDES does not have a common definition, it usually refers to a class of storage technologies with discharge duration that exceeds 6 hours [twitchell_defining_2023]. will be essential [cebulla_how_2018]. This need becomes particularly acute in the full or partial absence of alternative clean, firm resources, such as new nuclear power or natural‐gas plants equipped with carbon capture and storage (CCS) [de_sisternes_value_2016].
@@
Integration of LDES faces at least two challenges. First, many LDES technologies are capital-intensive and relatively immature. To guide technology development, a recent research stream explored cost thresholds for LDES to appear in least-cost capacity mixes, concluding that significant reductions in energy-capacity costs are needed to compete with other firm capacity resources [sepulveda_design_2021]. Second, the business models that allow LDES to valorize all its operational and environmental benefits for the system [cole_peaking_2023] have not yet been established. In energy markets with price caps lower than the value of lost load (VOLL), a business model purely based on energy arbitrage remains uneconomic, with narrow price spreads failing to yield sufficient returns to fully support investments [schmidt_monetizing_2023].
@@
Pure arbitrage is equally unattractive for short-duration energy storage (SDES), e.g., electrochemical batteries. However, in the case of SDES, stacking revenues from multiple market products such as energy and ancillary services [tian_stacked_2018], combined with generous subsidies, such as tax credits or grants [dougherty_improving_2021], has boosted batteries’ profits and deployment. These two drivers have created robust SDES business models, leading to a sixfold surge of global annual installations since 2020 [iea_batteries_2024]. However, these models will likely not be transferable to LDES. Short-term applications, such as frequency regulating reserves and diurnal arbitrage, will likely continue to be dominated by SDES and additionally face market saturation risks [pollitt_competition_2021]. Furthermore, subsidies and grants can become unsustainable if a large number of clean energy assets need support.
@@
One benefit of LDES lies in its peaking potential and its contribution to system adequacy [cole_peaking_2023]. As such, in decarbonized power systems, LDES primarily replaces dispatchable conventional peakers, limits renewable curtailment, and ensures adequacy during contingencies or prolonged periods with low‑VRE generation (so‑called ‘energy droughts’ [bracken_standardized_2024]). Consequently, as with peaking units, a substantial portion of the LDES revenues is likely to be realized during scarcity or near-scarcity events. In markets with price caps lower than the VOLL, insufficiently high price spikes during scarcity events could create a missing money problem [newbery_missing_2016], challenging the viability of investment in LDES. Furthermore, the concentration of scarcity events within particular weather years can amplify interannual profit variability, undermining investor confidence even when wholesale energy price arbitrage provides sufficient revenue over the financial lifetime of the asset [mallapragada_electricity_2023].
@@
Capacity remuneration mechanisms (CRMs), such as capacity markets (CMs), address missing money and revenue volatility in price‑capped markets. However, CMs were originally designed with thermal generation in mind, and their effectiveness in contexts with substantial ES capacity, particularly LDES, remains understudied. Recent research has focused on accurately valuing the contribution of energy storage to system adequacy. Zachary et al. [zachary_integration_2022] demonstrated that a change in expected unserved energy (EUE) is a suitable metric for evaluating the contribution of resources to system adequacy, extending the analyses to VRE and storage (yet with strong assumptions about its dispatch). They also demonstrated that a marginal, rather than an average, approach could yield accurate investment signals [^2] Marginal capacity credit reflects the incremental contribution of a new resource to system adequacy, while average capacity credit captures the overall contribution of all existing resources of that type.. Although marginal accreditation methods are more complex to apply, multiple CM operators, including NYISO [mohrman_capacity_2022], ISO-NE [zhao_resource_2022], PJM [noauthor_order_2024], and GB [national_energy_system_operator_storage_2024], have recently begun to progress in this direction.
@@
The shift towards marginal accreditation is promising, as investment decisions are highly sensitive to the chosen methodology and underlying assumptions [mertens_capacity_2021]. Yet, even when marginal credits are used for ES and VRE, there is no general proof ensuring that investments will be efficient. Wang et al. [wang_crediting_2022] demonstrated through simulations that marginal accreditation in CMs can provide price signals that encourage near-efficient investment for VRE and SDES. While these results are encouraging for the effectiveness of CMs in the near future, with narrow participation of ES, they are of limited use for future systems where LDES may play a dominant role.
@@
In general, most existing studies focus on ES with a single value (usually short) of duration or study systems where ES is a minor contributor to a system’s resource adequacy [wang_crediting_2022, opathella_novel_2019]. Many that include ES in CRMs often make simplified assumptions about accreditation [sanchez_jimenez_capacity_2024]. These simplifications make the results of existing studies less relevant for systems with LDES. The duration of LDES, for many emerging technologies, is a design choice independent of the power capacity. Moreover, LDES duration can range from a few to several hundred hours. Different durations yield distinct contributions to system adequacy, complicating decisions for both investors and regulators.
@@
We contribute to the literature in several ways. Methodologically, we extend the framework of [wang_crediting_2022] by incorporating a price-elastic demand to address degeneracy and improve price interpretability under high VRE and storage penetration. This is also the first study to endogenise continuous ES accreditation within an investment model, using piecewise linear duration functions. Unlike previous studies (e.g., [mertens_capacity_2021], [askeland_equilibrium_2019]) that estimate capacity credits for a discrete set of predefined duration levels, the proposed continuous ES accreditation curves provide capacity credits for a continuous set of durations, allowing the model to choose optimal storage configurations.
Our analysis offers new insights into market-based adequacy in low-carbon systems with high LDES levels. We confirm that the missing money problem remains significant under price caps, but declines as storage is able to capture wider spreads by avoiding VRE curtailment. We also show that the price cap affects revenues and operations below the cap itself, particularly in cases of low CO2 emissions. Well-calibrated CMs with marginal EUE-based accreditation maintain near-efficient investment signals, but their effectiveness is lower under lower decarbonization targets.
These findings are relevant for CM design and resource adequacy methodologies. They also inform policymakers considering LDES support mechanisms, such as the UK’s cap-and-floor scheme [department_for_energy_security__net_zero_long_2024] by highlighting the interplay between accreditation, revenue volatility, and revenue sufficiency.
@@ II-A1 Generators
Let $\mathcal{G}$ be a set of electricity generators. Each generator has three decision variables: its installed capacity ($c_{g}$), the capacity contracted by the CM ($c^{\text{CM}}_{g}$), and the electricity produced in each time interval ($q_{\omega tg}$), which are in the set $\Gamma^{g}=\left\{q_{\omega tg},c_{g},c^{\text{CM}}_{g}\geq 0\right\}$. Acting as a price taker, each generator solves the optimization problem (1) below.
$\forall g\in\mathcal{G}\quad\max_{\Gamma^{g}}\sum_{\omega t}w_{\omega t}\delta_{\omega}\left(\lambda^{E}_{\omega t}-C_{g}^{V}-\lambda^{\text{CT}}\text{EF}_{g}\right)q_{\omega tg}$
$+c^{\text{CM}}_{g}\lambda^{\text{CM}}-c_{g}\left(\widetilde{I}_{g}+I_{g}^{F}\right)\hskip 0.0pt$ (1a)
$q_{\omega tg}\leq c_{g}A_{\omega tg}\hskip 0.0pt\quad\forall\omega,t;$ (1b)
$c^{\text{CM}}_{g}\leq\text{CC}_{g}c_{g}\hskip 0.0pt$ (1c)
@@ II-A2 Storage
Let set $\mathcal{S}$ include all ES market participants. ES participants determine power and energy capacity independently. Additionally, ES decision variables include the capacity contracted by the CM, its initial state-of-charge (SOC), its charging, and discharging. All decision variables are in $\Gamma_{s}=\left\{q_{\omega,t,s}^{\text{dis}},q_{\omega,t,s}^{ch},c_{s}^{P},c_{s}^{E},e_{\omega,t,s},e^{\text{init}},c^{\text{CM}}_{s}\geq 0\right\}$ and their values are determined by solving the optimization problem (2).
@@
$+c^{\text{CM}}_{s}\lambda^{\text{CM}}-c_{s}^{P}\left({\widetilde{I}}_{s}^{P}+I_{s}^{P.F}\right)-c_{s}^{E}\left({\widetilde{I}}_{s}^{E}+I_{s}^{E.F}\right)\hskip 0.0pt$ (2a)
@@
$e_{\omega ts}=e_{s}^{\text{init}}+w_{\omega t}(q_{\omega ts}^{\text{ch}}\eta_{s}^{\text{ch}}-\frac{q_{\omega ts}^{\text{dis}}}{\eta_{s}^{\text{dis}}})\quad\forall\;\omega,\;t=1$ (2e)
$e_{\omega ts}\geq e_{s}^{\text{init}}\hskip 0.0pt\;\forall\;\omega,\;t=T$ (2f)
@@
$c^{\text{CM}}_{s}\leq c_{s}^{P}f^{\text{CC}}(c_{s}^{E},c_{s}^{P})\hskip 0.0pt$ (2h)
Eq. (2a) is the ES’s objective function, maximising its profit. Eq. (2b)-(2c) limit charging and discharging based on installed power capacity. Eq. (2d)-(2e) calculate the energy stored at the end of a time interval of duration $w_{\omega t}$ by considering the energy stored at the start of the interval and the charging, and discharging during the interval. Eq. (2f)-(2g) limit the energy stored relative to the energy stored at the start of the model’s horizon and the installed energy capacity, respectively [^3] Constraining end-of-horizon stored energy to exceed initial levels may be restrictive, but it provides a practical way to handle model end effects under limited foresight. It is also consistent with EU rules requiring minimum gas storage levels before winter [european_comission_amending_2025]..
Finally, (2h) limits the capacity contracted by the CM, $c^{\text{CM}}_{s}$, to the qualifying capacity of ES. We determine the capacity credit of ES in a novel manner as a function, $f^{\text{CC}}$, of ES duration $\zeta_{s}$ $\eta_{s}^{\text{dis}}c_{s}^{E}/c_{s}^{P}$. This new way allows us to properly consider the capacity market contributions of a range of ES durations. The function $f^{CC}$ is piece-wise linear with $|L|$ segments, where each segment $l$ covers a range of durations $\zeta_{sl-1},\zeta_{sl}]$. As such, $f^{\text{CC}}$ is represented as
$f^{\text{CC}}(\zeta_{s})=\alpha_{ls}+\beta_{ls}\zeta_{s},\quad\forall\;\zeta_{s}\in[\zeta_{sl-1},\zeta_{sl}],l\in\mathcal{L}$ (3)
Capacity credit curves are monotonically increasing and concave with respect to duration [paul_denholm_beyond_2023]. In other words, the marginal contribution to adequacy diminishes as duration increases (i.e., the value of $\beta_{ls}$ is positive but lower for higher durations). We estimate the coefficients of $f^{\text{CC}}$, i.e., $\alpha_{ls},\beta_{ls}$, using piece-wise linear regression.
In the maximisation problem (2), constraint (2h) can be included as is because $f^{\text{CC}}$ is concave. If $f^{\text{CC}}$ was not concave, binary variables and associated constraints would be needed to ensure that the segments of the piecewise curve are activated in a logical order. Multiplying both sides of (3) with the installed power capacity $c_{s}^{P}$, (2h) becomes:
$c^{\text{CM}}_{s}\leq\alpha_{ls}c_{s}^{P}+\beta_{ls}\eta_{s}^{\text{dis}}c_{s}^{E},\quad\forall l\in\mathcal{L}$ (4)
@@ II-A3 Consumers
We assume a single representative agent acting on behalf of all consumers. Its decision variables include the electricity consumption and capacity purchased in the CM, grouped in the set $\Gamma^{\text{CM}}=\left\{\left.d^{E}_{\omega t}\right|t\in\mathcal{T},\omega\in\Omega,d^{\text{CM}}\geq 0\right\}$. Consumers maximise the expected consumer surplus, which is defined as the difference between the integral of willingness to pay (WTP) functions for energy $f^{E}$ and capacity $f^{\text{CM}}$ and the procurement cost (product of price and quantity purchased).
$\max_{\Gamma^{C}}\sum_{\omega t}w_{\omega t}\delta_{\omega}\Bigg[\int_{0}^{d^{E}_{\omega t}}f^{E}_{\omega t}(x)dx-\lambda^{E}_{\omega t}d^{E}_{\omega t}\Bigg]\\ +\int_{0}^{d^{\text{CM}}}f^{\text{CM}}(x)dx-\lambda^{\text{CM}}d^{\text{CM}}\hskip 0.0pt$ (5a)
The demand curves for electricity can take diverse shapes, ranging from perfectly price-inelastic to perfectly price-elastic curves. In this work, in line with [de_jonghe_optimal_2012], we assume that the demand curve for electricity at each time interval has two segments. The first segment is perfectly elastic (horizontal) for the willingness to pay for electricity up to the quantity demanded of $D^{\text{fix}}_{\omega t}$ is constant at $B$. The second segment is a price-elastic linear demand curve, with willingness to pay ranging from 0 to $B$ for consumption between 0 and $D^{\text{flex}}_{\omega t}$. This demand curve helps ensure price uniqueness and resolve dual degeneracy issues in systems with high shares of renewables and storage. The WTP function for energy $f^{E}$ is included in the model using the equations below.
@@
$d_{\omega t}^{\text{flex}}\leq D_{\omega t}^{\text{flex}}\quad\forall\omega,t\hskip 0.0pt$ (6c)
The CM function is represented similarly and often has more segments with varying levels of elasticity. The CM function $f^{\text{CM}}(d^{\text{CM}})$ represents the WTP for capacity, and is a monotonically decreasing function. It consists of $N$ segments, with differing levels of elasticity. The first segment is perfectly price-elastic (horizontal), willing to pay up to the capacity market price cap $B^{C}_{1}$ for any capacity between 0 and $D^{\text{C-fix}}$. In the last segment, the WTP $B^{C}_{N}$ decreases to zero as $d^{C}$ reaches the maximum capacity the CM aims to procure. The N2N-2 segments in between are flexible (lengths $D_{n}^{\text{C-flex}}$, starting prices $B^{C}_{n}$, with $B^{C}_{N}=0$) and they are parametrized based on the net cost of new entry (CONE). Given this structure, the function can be represented as:
$\int_{0}^{d^{\text{CM}}=d^{\text{fix}}+\sum_{n=1}^{N-1}d_{n}^{\text{flex}}}f^{\text{CM}}(x)\,dx=B^{C}_{1}d^{\text{C-fix}}+\\ \sum_{n=1}^{N-1}B^{C}_{n}d_{n}^{\text{C-flex}}-\sum_{n=1}^{N-1}\frac{(d_{n}^{\text{C-flex}})^{2}(B^{C}_{n}-B^{C}_{n+1})}{2D_{n}^{\text{C-flex}}},\hskip 0.0pt$ (7)